OpenStudy (anonymous):

-53 + n = -28 please show work?

OpenStudy (math&ing001):

If you're looking to find n, add 53 on both sides : -53 + n = -28 +53 +53 n = -28 +53

OpenStudy (anonymous):

n = 25

OpenStudy (zbrandz23):

Your gonna have to add 28 on both sides first, then subtract and you'll get n by itself.